# Jacobian Descent For Multi-Objective Optimization

### v1

@@ B.7 CAGrad
$\mathcal{A}_{\text{CAGrad}}$ is described in [12]. It is parameterized by $c\in\left[0,1\right[{}$. If $c=0$, this is equivalent to $\mathcal{A}_{\text{Mean}}$. Therefore, we restrict our analysis to the case $c>0$. For any $J\in\mathbb{R}^{m\times n}$, let $\bar{\boldsymbol{g}}$ be the average gradient $\frac{1}{m}J^{\top}\cdot{\boldsymbol{1}}$. The aggregated vector is defined as the minimizer $d\in{\mathbb{R}}^{n}$ of the following optimization problem
@@
$\operatorname*{arg\,max}_{\begin{subarray}{c}{\boldsymbol{d}}\in{\mathbb{R}}^{n}:\\ \|{\boldsymbol{d}}-\overline{\boldsymbol{g}}\|\varleq c\|\overline{\boldsymbol{g}}\|\end{subarray}}\min_{i\in[m]}{\boldsymbol{e}}_{i}^{\top}J{\boldsymbol{d}}$
@@
For $10<c<$ 1, $\mathcal{A}_{\text{CAGrad}}$ can have conflict. If $J=\begin{bmatrix}2&0\\ -2a-2&2\end{bmatrix}$, then $\overline{\boldsymbol{g}}=\begin{bmatrix}-a&1\end{bmatrix}^{\top}$ and $\|\overline{\boldsymbol{g}}\|=\sqrt{a^{2}+1}$. Observe that any ${\boldsymbol{d}}\in{\mathbb{R}}^{n}$ satisfying the constraint $\|{\boldsymbol{d}}-\overline{\boldsymbol{g}}\|\varleq c\|\overline{\boldsymbol{g}}\|$ has first coordinate at most $-a+c\sqrt{a^{2}+1}$. Suppose that $a$ is such that $-a+c\sqrt{a^{2}+1}<0$, i.e. $\sqrt{\frac{c^{2}}{1-c^{2}}}<a$. For such an $a$, any feasible $d$ has a negative first coordinate, making it conflict with the first row of $J$. $\mathcal{A}_{\text{CAGrad}}$ is thus not non-conflicting.
@@
Note that if we generalize to the case $c\vargeq 1$, as suggested in the original paper, then ${\boldsymbol{d}}={\boldsymbol{0}}$ becomes feasible, which yields $\min_{i\in[m]}{\boldsymbol{e}}_{i}^{\top}J{\boldsymbol{d}}=0$. Therefore the optimal ${\boldsymbol{d}}$ is such that $0\varleq\min_{i\in[m]}{\boldsymbol{e}}_{i}^{\top}J{\boldsymbol{d}}$, i.e. ${\boldsymbol{0}}\preceq J{\boldsymbol{d}}$. With $c\vargeq 1$, $\mathcal{A}_{\text{CAGrad}}$ would thus be non-conflicting.
In [12], they formulate $\mathcal{A}_{\text{CAGrad}}$ using its dual formulation. If ${\boldsymbol{w}}\in{\mathbb{R}}^{m}$ is a solution to
$\min_{\begin{subarray}{c}{\boldsymbol{0}}\preceq{\boldsymbol{w}}:\\ {\boldsymbol{1}}^{\top}{\boldsymbol{w}}=1\end{subarray}}{\boldsymbol{1}}^{\top}JJ^{\top}{\boldsymbol{w}}+c\cdot\|J^{\top}{\boldsymbol{1}}\|\cdot\|J^{\top}{\boldsymbol{w}}\|$
then the aggregation is $\frac{1}{m}J^{\top}\left({\boldsymbol{1}}+\frac{c\|J^{\top}{\boldsymbol{1}}\|}{\|J^{\top}{\boldsymbol{w}}\|}{\boldsymbol{w}}\right)$. Therefore, $\mathcal{A}_{\text{CAGrad}}$ is weighted.
@@
$\mathcal{A}_{\text{CAGrad}}$ is not linear under scaling. We sketch a proof of this. Suppose that $J=\begin{bmatrix}2&0\\ 0&2a\end{bmatrix}$. Note that $\overline{\boldsymbol{g}}=\begin{bmatrix}1&a\end{bmatrix}^{\top}$ and $\|\overline{\boldsymbol{g}}\|=\sqrt{1+a^{2}}$. One can show that the constraint $\|{\boldsymbol{d}}-\overline{\boldsymbol{g}}\|\varleq c\|\overline{\boldsymbol{g}}\|$ needs to be satisfied with equality since, otherwise, we can scale ${\boldsymbol{d}}$ to make the objective larger. Substituting $J$ in $\min_{i\in[m]}{\boldsymbol{e}}_{i}^{\top}J{\boldsymbol{d}}$ yields $2\min(d_{1},ad_{2})$. For any $a$ such that $c\sqrt{1+a^{2}}+1<a^{2}$, it can be shown that the optimal ${\boldsymbol{d}}$ is such that $d_{1}<ad_{2}$. In that case the inner minimum over $i$ is $2d_{1}$ and, to satisfy $\|{\boldsymbol{d}}-\overline{\boldsymbol{g}}\|=c\|\overline{\boldsymbol{g}}\|$, the KKT conditions over the Lagrangian yield ${\boldsymbol{d}}-\overline{\boldsymbol{g}}~{}\propto~{}\nabla_{\boldsymbol{d}}d_{1}=\begin{bmatrix}1&0\end{bmatrix}^{\top}$. This yields ${\boldsymbol{d}}=\begin{bmatrix}c\cdot\|\overline{\boldsymbol{g}}\|+1\\ a\end{bmatrix}=\begin{bmatrix}c\sqrt{1+a^{2}}+1\\ a\end{bmatrix}$. This is not affine in $a$; therefore, $\mathcal{A}_{\text{CAGrad}}$ is not linear under scaling.
@@ B.8 RGW
$\mathcal{A}_{\text{RGW}}$ is defined in [14] as the weighted sum of the rows of the input matrix, with a random weighting. The weighting is obtained by sampling $m$ i.i.d. normally distributed random variables and applying a softmax. Formally, $\mathcal{A}_{\text{RGW}}(J)=J^{\top}\cdot\sigma(W)$, with $W\sim N({\boldsymbol{0}},I)$. By design, $\mathcal{A}_{\text{RGW}}$ is thus weighted.
When $J=\begin{bmatrix}1\\ -1\end{bmatrix}$, the only non-conflicting solution is ${\boldsymbol{0}}$. However, $\mathds{P}[\mathcal{A}_{\text{RGW}}(J)={\boldsymbol{0}}]=0$. $\mathcal{A}_{\text{RGW}}$ is thus not non-conflicting.
$\mathcal{A}_{\text{RGW}}(\operatorname*{diag}({\boldsymbol{c}})\cdot J)=(\operatorname*{diag}({\boldsymbol{c}})\cdot J)^{\top}\cdot\sigma(W)=J^{\top}\cdot\operatorname*{diag}({\boldsymbol{c}})\cdot\sigma(W)=J^{\top}\cdot\operatorname*{diag}(\sigma(W))\cdot{\boldsymbol{c}}$. Therefore, $\mathcal{A}_{\text{RGW}}$ is linear under scaling in distribution.
@@ B.9 Nash-MTL
Nash-MTL is described in [15]. Unfortunately, we were not able to verify the proof of Claim 3.1, and we believe that the official implementation of Nash-MTL may mismatch the desired objective by which it is defined. Therefore, we only analyze the initial objective even though our experiments for this aggregator are conducted with the official implementation.
Let $J\in\mathbb{R}^{m\times n}$ and $\varepsilon>0$. Let also $B_{\varepsilon}=\{{\boldsymbol{d}}\in{\mathbb{R}}^{n}:\|{\boldsymbol{d}}\|\varleq\varepsilon,{\boldsymbol{0}}\preceq J{\boldsymbol{d}}\}$. $\mathcal{A}_{\text{Nash-MTL}}$ is then defined as
$\mathcal{A}_{\text{Nash-MTL}}(J)$ $\in\operatorname*{arg\,max}_{{\boldsymbol{d}}\in B_{\varepsilon}}\sum_{i=1}^{m}\log({\boldsymbol{e}}_{i}^{\top}J{\boldsymbol{d}})$
By the constraint, $\mathcal{A}_{\text{Nash-MTL}}$ is non-conflicting. If an aggregator $\mathcal{A}$ is linear under scaling, it should be the case that $\mathcal{A}(aJ)=a\mathcal{A}(J)$ for any scalar $a>0$ and any $J\in\mathbb{R}^{m\times n}$. However, $\log(a{\boldsymbol{e}}_{i}^{\top}J{\boldsymbol{d}})=\log({\boldsymbol{e}}_{i}^{\top}J{\boldsymbol{d}})+\log(a)$. This means that scaling by a scalar does not impact aggregation. Since this is not the trivial ${\boldsymbol{0}}$ aggregator, $\mathcal{A}_{\text{Nash-MTL}}$ is not linear under scaling.
If ${\boldsymbol{d}}^{\prime}$ is the projection of ${\boldsymbol{d}}$ onto the span of the rows of $J$, then $\|{\boldsymbol{d}}^{\prime}\|\varleq\|{\boldsymbol{d}}\|$ and $J{\boldsymbol{d}}=J{\boldsymbol{d}}^{\prime}$. Therefore, ${\boldsymbol{d}}^{\prime}$ is feasible and has the same associated objective value. Therefore, without loss of generality, $\mathcal{A}_{\text{Nash-MTL}}$ is weighted.
@@ B.10 Aligned-MTL
The Aligned-MTL method for balancing the Jacobian is described in [16]. For simplicity, we fix the vector of preferences to $\frac{1}{m}{\boldsymbol{1}}$, but the proofs can be adapted for any non-trivial vector. Given $J\in\mathbb{R}^{m\times n}$, compute the eigen-decomposition of $G=JJ^{\top}=V\Sigma^{2}V^{\top}$. Let $\Sigma^{\dagger}$ be the diagonal matrix whose non-zero elements are the inverse of corresponding non-zero diagonal elements of $\Sigma$. Let $\sigma_{\min}=\min_{i\in[m],\Sigma_{ii}\neq 0}\Sigma_{ii}$. Then the aggregation by $\mathcal{A}_{\text{Aligned-MTL}}$ is equal to $\frac{1}{m}J^{\top}\cdot{\boldsymbol{w}}$, with ${\boldsymbol{w}}=\sigma_{\min}\cdot V\Sigma^{\dagger}V^{\top}\cdot{\boldsymbol{1}}$. It follows immediately that $\mathcal{A}_{\text{Aligned-MTL}}$ is weighted.
If the SVD of $J$ is $V\Sigma U^{\top}$, then $J^{\top}{\boldsymbol{w}}=\sigma_{\min}UPV^{\top}{\boldsymbol{1}}$ with $P=\Sigma^{\dagger}\Sigma$ a diagonal projection matrix with $1$s corresponding to non zero elements of $\Sigma$ and 00s everywhere else. Further, $J\cdot\mathcal{A}_{\text{Aligned-MTL}}(J)=\frac{\sigma_{\min}}{m}\cdot V\Sigma V^{\top}{\boldsymbol{1}}$. Taking $V=\frac{1}{2}\begin{bmatrix}\sqrt{3}&1\\ -1&\sqrt{3}\end{bmatrix}$, and $\Sigma=\begin{bmatrix}1&0\\ 0&0\end{bmatrix}$ yields $J\cdot\mathcal{A}_{\text{Aligned-MTL}}(J)=\frac{1}{2}\cdot V\Sigma V^{\top}{\boldsymbol{1}}=\frac{1}{8}\begin{bmatrix}-3+\sqrt{3}&1-\sqrt{3}\end{bmatrix}^{\top}$ which is not non-negative. $\mathcal{A}_{\text{Aligned-MTL}}$ is thus not non-conflicting.
If $J=\begin{bmatrix}1&0\\ 0&a\end{bmatrix}$, then $U=V=I$, $\Sigma=J$. For $0<a\varleq 1$, $\sigma_{\min}=a$, therefore $\mathcal{A}_{\text{Aligned-MTL}}(J)=\frac{a}{2}\cdot{\boldsymbol{1}}$. For $1<a$, $\sigma_{\min}=1$ and therefore $\mathcal{A}_{\text{Aligned-MTL}}(J)=\frac{{\boldsymbol{1}}}{2}$, which makes $\mathcal{A}_{\text{Aligned-MTL}}$ not linear under scaling.

### v2

@@ B.7 CAGrad
$\mathcal{A}_{\text{CAGrad}}$ is described in Liu et al. (2021a). It is parameterized by $c\in\left[0,1\right[{}$. If $c=0$, this is equivalent to $\mathcal{A}_{\text{Mean}}$. Therefore, we restrict our analysis to the case $c>0$. For any $J\in\mathbb{R}^{m\times n}$, let $\bar{\bm{g}}$ be the average gradient $\frac{1}{m}J^{\top}\cdot{\bm{1}}$, and let ${\bm{e}_{i}}^{\top}J$ denote the $i$’th row of $J$. The aggregation is then defined as
@@
$\mathcal{A}_{\text{CAGrad}}(J)\in\operatorname*{arg\,max}_{\begin{subarray}{c}{\bm{d}}\in{\mathbb{R}}^{n}:\\ \|{\bm{d}}-\overline{\bm{g}}\|\varleq c\|\overline{\bm{g}}\|\end{subarray}}\min_{i\in[m]}{\bm{e}_{i}}^{\top}J{\bm{d}}$ (31)
  Non-conflicting.
@@
Let $J=\begin{bmatrix}2&0\\ -2a-2&2\end{bmatrix}$, with $a$ satisfying $-a+c\sqrt{a^{2}+1}<0$. We have $\overline{\bm{g}}=\begin{bmatrix}-a&1\end{bmatrix}^{\top}$ and $\|\overline{\bm{g}}\|=\sqrt{a^{2}+1}$. Observe that any ${\bm{d}}\in{\mathbb{R}}^{n}$ satisfying the constraint $\|{\bm{d}}-\overline{\bm{g}}\|\varleq c\|\overline{\bm{g}}\|$ has first coordinate at most $-a+c\sqrt{a^{2}+1}$. Because $-a+c\sqrt{a^{2}+1}<0$, any feasible $d$ has a negative first coordinate, making $d$ conflict with the first row of $J$. For any $c\in[0,1[$, $-a+c\sqrt{a^{2}+1}<0$ is equivalent to $\sqrt{\frac{c^{2}}{1-c^{2}}}<a$. Thus, this provides a counter-example to the non-conflicting property for any $c\in[0,1[$, i.e. $\mathcal{A}_{\text{CAGrad}}$ is not non-conflicting.
@@
If we generalize to the case $c\vargeq 1$, as suggested in the original paper, then ${\bm{d}}={\bm{0}}$ becomes feasible, which yields $\min_{i\in[m]}{\bm{e}_{i}}^{\top}J{\bm{d}}=0$. Therefore the optimal ${\bm{d}}$ satisfies $0\varleq\min_{i\in[m]}{\bm{e}_{i}}^{\top}J{\bm{d}}$, i.e. ${\bm{0}}\preceq J{\bm{d}}$. With $c\vargeq 1$, $\mathcal{A}_{\text{CAGrad}}$ would thus be non-conflicting.
  Linear under scaling (sketch of proof).
@@
Let $J=\begin{bmatrix}2&0\\ 0&2a\end{bmatrix}$, then $\overline{\bm{g}}=\begin{bmatrix}1&a\end{bmatrix}^{\top}$ and $\|\overline{\bm{g}}\|=\sqrt{1+a^{2}}$. One can show that the constraint $\|{\bm{d}}-\overline{\bm{g}}\|\varleq c\|\overline{\bm{g}}\|$ needs to be satisfied with equality since, otherwise, we can scale ${\bm{d}}$ to make the objective larger. Substituting $J$ in $\min_{i\in[m]}{\bm{e}_{i}}^{\top}J{\bm{d}}$ yields $2\min(d_{1},ad_{2})$. For any $a$ satisfying $c\sqrt{1+a^{2}}+1<a^{2}$, it can be shown that the optimal ${\bm{d}}$ satisfies $d_{1}<ad_{2}$. In that case the inner minimum over $i$ is $2d_{1}$ and, to satisfy $\|{\bm{d}}-\overline{\bm{g}}\|=c\|\overline{\bm{g}}\|$, the KKT conditions over the Lagrangian yield ${\bm{d}}-\overline{\bm{g}}~{}\propto~{}\nabla_{\bm{d}}d_{1}=\begin{bmatrix}1&0\end{bmatrix}^{\top}$. This yields ${\bm{d}}=\begin{bmatrix}c\cdot\|\overline{\bm{g}}\|+1\\ a\end{bmatrix}=\begin{bmatrix}c\sqrt{1+a^{2}}+1\\ a\end{bmatrix}$. This is not affine in $a$; therefore, $\mathcal{A}_{\text{CAGrad}}$ is not linear under scaling.
  Weighted.
In Liu et al. (2021a), $\mathcal{A}_{\text{CAGrad}}$ is formulated via its dual: $\mathcal{A}_{\text{CAGrad}}(J)=\frac{1}{m}J^{\top}\left({\bm{1}}+\frac{c\|J^{\top}{\bm{1}}\|}{\|J^{\top}{\bm{w}}\|}{\bm{w}}\right)$, with ${\bm{w}}\in\operatorname*{arg\,min}_{\begin{subarray}{c}{\bm{w}}\in\Delta(m)\end{subarray}}{\bm{1}}^{\top}JJ^{\top}{\bm{w}}+c\cdot\|J^{\top}{\bm{1}}\|\cdot\|J^{\top}{\bm{w}}\|$, where is $\Delta(m)$ the probability simplex of dimension $m$. Therefore, $\mathcal{A}_{\text{CAGrad}}$ is weighted.
@@ B.8 RGW
$\mathcal{A}_{\text{RGW}}$ is defined in Lin et al. (2021) as the weighted sum of the rows of the input matrix, with a random weighting. The weighting is obtained by sampling $m$ i.i.d. normally distributed random variables and applying a softmax. Formally,
$\mathcal{A}_{\text{RGW}}(J)$ $=J^{\top}\cdot\mathrm{softmax}({\mathbf{w}})$ (32)
$\text{with}\hskip 12.80365pt{\mathbf{w}}$ $\sim\mathcal{N}({\bm{0}},I)$ (33)
  Non-conflicting.
When $J=\begin{bmatrix}1\\ -2\end{bmatrix}$, the only non-conflicting solution is ${\bm{0}}$. However, $\mathds{P}[\mathcal{A}_{\text{RGW}}(J)={\bm{0}}]=0$, i.e. $\mathcal{A}_{\text{RGW}}(J)$ is in the dual cone of the rows of $J$ with probability 00. $\mathcal{A}_{\text{RGW}}$ is thus not non-conflicting. Here,
[…]
$=-\mathbb{E}\left[\frac{e^{{\textnormal{w}}_{1}}}{e^{{\textnormal{w}}_{1}}+e^{{\textnormal{w}}_{2}}}\right]$ $\left(\begin{subarray}{c}{\mathbf{w}}\sim\mathcal{N}({\bm{0}},I)\end{subarray}\right)$
$<0$ $\left(\begin{subarray}{c}0\prec e^{{\mathbf{w}}}\end{subarray}\right)$
so $\mathcal{A}_{\text{RGW}}$ is neither non-conflicting in expectation.
  Linear under scaling.
$\mathcal{A}_{\text{RGW}}\big{(}\operatorname*{diag}({\bm{c}})\cdot J\big{)}=\big{(}\operatorname*{diag}({\bm{c}})\cdot J\big{)}^{\top}\cdot\mathrm{softmax}({\mathbf{w}})=J^{\top}\cdot\operatorname*{diag}({\bm{c}})\cdot\mathrm{softmax}({\mathbf{w}})=J^{\top}\cdot\operatorname*{diag}\big{(}\mathrm{softmax}({\mathbf{w}})\big{)}\cdot{\bm{c}}$. We thus have $\mathcal{A}_{\text{RGW}}\big{(}\operatorname*{diag}({\bm{c}})\cdot J\big{)}=\mathfrak{J}^{\top}\cdot{\bm{c}}$ with $\mathfrak{J}^{\top}=J^{\top}\cdot\operatorname*{diag}\big{(}\mathrm{softmax}({\mathbf{w}})\big{)}$. Therefore, $\mathcal{A}_{\text{RGW}}$ is linear under scaling.
  Weighted.
By (32), $\mathcal{A}_{\text{RGW}}$ is weighted.
@@ B.9 Nash-MTL
Nash-MTL is described in Navon et al. (2022). Unfortunately, we were not able to verify the proof of Claim 3.1, and we believe that the official implementation of Nash-MTL may mismatch the desired objective by which it is defined. Therefore, we only analyze the initial objective even though our experiments for this aggregator are conducted with the official implementation.
Let $J\in\mathbb{R}^{m\times n}$ and $\varepsilon>0$. Let also $B_{\varepsilon}=\big{\{}{\bm{d}}\in{\mathbb{R}}^{n}:\|{\bm{d}}\|\varleq\varepsilon,{\bm{0}}\preceq J{\bm{d}}\big{\}}$. With ${\bm{e}_{i}}^{\top}J$ denoting the $i$’th row of $J$, $\mathcal{A}_{\text{Nash-MTL}}$ is then defined as
$\mathcal{A}_{\text{Nash-MTL}}(J)=\operatorname*{arg\,max}_{{\bm{d}}\in B_{\varepsilon}}\sum_{i\in[m]}\log\big{(}{\bm{e}_{i}}^{\top}J{\bm{d}}\big{)}$ (34)
  Non-conflicting.
By the constraint, $\mathcal{A}_{\text{Nash-MTL}}$ is non-conflicting.
  Linear under scaling.
If an aggregator $\mathcal{A}$ is linear under scaling, it should be the case that $\mathcal{A}(aJ)=a\mathcal{A}(J)$ for any scalar $a>0$ and any $J\in\mathbb{R}^{m\times n}$. However, $\log(a{\bm{e}_{i}}^{\top}J{\bm{d}})=\log({\bm{e}_{i}}^{\top}J{\bm{d}})+\log(a)$. This means that scaling by a scalar does not impact aggregation. Since this is not the trivial ${\bm{0}}$ aggregator, $\mathcal{A}_{\text{Nash-MTL}}$ is not linear under scaling.
  Weighted.
Suppose towards contradiction that ${\bm{d}}$ is both optimal for (34) and not in the span of $J^{\top}$. Let ${\bm{d}}^{\prime}$ be the projection of ${\bm{d}}$ onto the span of $J^{\top}$. Since $\|{\bm{d}}^{\prime}\|<\|{\bm{d}}\|<\varepsilon$ and $J{\bm{d}}=J{\bm{d}}^{\prime}$, we have $J{\bm{d}}\prec J\left(\frac{\|{\bm{d}}\|}{\|{\bm{d}}^{\prime}\|}{\bm{d}}^{\prime}\right)$, contradicting the optimality of ${\bm{d}}$. Therefore, $\mathcal{A}_{\text{Nash-MTL}}$ is weighted.
@@ B.10 Aligned-MTL
The Aligned-MTL method for balancing the Jacobian is described in Senushkin et al. (2023). For simplicity, we fix the vector of preferences to $\frac{1}{m}{\bm{1}}$, but the proofs can be adapted for any non-trivial vector. Given $J\in\mathbb{R}^{m\times n}$, let $V\Sigma^{2}V^{\top}$ be the eigen-decomposition of $JJ^{\top}$, let $\Sigma^{\dagger}$ be the diagonal matrix whose non-zero elements are the inverse of corresponding non-zero diagonal elements of $\Sigma$ and let $\sigma_{\min}=\min_{i\in[m],\Sigma_{ii}\neq 0}\Sigma_{ii}$. The aggregation is then defined as
$\mathcal{A}_{\text{Aligned-MTL}}(J)$ $=\frac{1}{m}J^{\top}\cdot{\bm{w}}$ (35)
$\text{with}\hskip 35.56593pt{\bm{w}}$ $=\sigma_{\min}\cdot V\Sigma^{\dagger}V^{\top}\cdot{\bm{1}}$ (36)
  Non-conflicting.
If the SVD of $J$ is $V\Sigma U^{\top}$, then $J^{\top}{\bm{w}}=\sigma_{\min}UPV^{\top}{\bm{1}}$ with $P=\Sigma^{\dagger}\Sigma$ a diagonal projection matrix with $1$s corresponding to non zero elements of $\Sigma$ and 00s everywhere else. Further, $J\cdot\mathcal{A}_{\text{Aligned-MTL}}(J)=\frac{\sigma_{\min}}{m}\cdot V\Sigma V^{\top}{\bm{1}}$. If $V=\frac{1}{2}\begin{bmatrix}\sqrt{3}&1\\ -1&\sqrt{3}\end{bmatrix}$ and $\Sigma=\begin{bmatrix}1&0\\ 0&0\end{bmatrix}$, we have $J\cdot\mathcal{A}_{\text{Aligned-MTL}}(J)=\frac{1}{2}\cdot V\Sigma V^{\top}{\bm{1}}=\frac{1}{8}\begin{bmatrix}3-\sqrt{3}&1-\sqrt{3}\end{bmatrix}^{\top}$ which is not non-negative. $\mathcal{A}_{\text{Aligned-MTL}}$ is thus not non-conflicting.
  Linear under scaling.
If $J=\begin{bmatrix}1&0\\ 0&a\end{bmatrix}$, then $U=V=I$, $\Sigma=J$. For $0<a\varleq 1$, $\sigma_{\min}=a$, therefore $\mathcal{A}_{\text{Aligned-MTL}}(J)=\frac{a}{2}\cdot{\bm{1}}$. For $1<a$, $\sigma_{\min}=1$ and therefore $\mathcal{A}_{\text{Aligned-MTL}}(J)=\frac{1}{2}\cdot{\bm{1}}$, which makes $\mathcal{A}_{\text{Aligned-MTL}}$ not linear under scaling.
  Weighted.
By (35), $\mathcal{A}_{\text{Aligned-MTL}}$ is weighted.